# Small Hurewicz and Menger sets which have large continuous images

### v1

@@ Abstract.
The research of the first and second authors was funded by the National Science Center, Poland Weave-UNISONO call in the Weave programme Project: Set-theoretic aspects of topological selections 2021/03/Y/ST1/00122. The research of the third author was funded in whole by the Austrian Science Fund (FWF) [I 5930].
@@ 1. Introduction
Throughout this paper, we assume that a set $X$ with a given topological property is a subset of the Cantor space ${2^{\omega}}$ or its square. A set $X$ is Menger if for any sequence $\mathcal{U}_{0},\allowbreak\mathcal{U}_{1},\allowbreak\dotsc$ of open covers of $X$, there are finite sets $\mathcal{F}_{0}\subseteq\mathcal{U}_{0},\mathcal{F}_{1}\subseteq\mathcal{U}_{1},\dotsc$ such that the family $\{\,\bigcup\mathcal{F}_{n}:n\in\omega\,\}$ covers $X$. If in addition the family $\{\,\bigcup\mathcal{F}_{n}:n\in\omega\,\}$ is a $\gamma$-cover of $X$, i.e., for every $x\in X$, the set $\{\,n:x\in\bigcup\mathcal{F}_{n}\,\}$ is co-finite, then the set $X$ is Hurewicz. Each $\sigma$-compact set is Hurewicz and each Hurewicz set is Menger. By the results of Bartoszyński–Shelah [5] and Tsaban–Zdomskyy [25], there are uniform constructions of sets in ZFC which show that none of these implications are reversible in the realm of sets of reals. The above properties are central in the combinatorial covering propeties theory and found applications in such areas as forcing [6], function spaces [20], Ramsey theory in algebra [24], combinatorics of discrete subspaces [1], hyperspaces with the Vietoris topology [9], products of Lindelöf spaces [2, 22] and products of paracompact spaces [21].
Each of the considered properties has a combinatorial characterization using continuous mappings into the Baire space ${\omega^{\omega}}$. A set $X$ is Menger (Hurewicz) if and only if every continuous image of $X$ into ${\omega^{\omega}}$ is non-dominating (bounded) [18]. It follows that each of these properties has a critical cardinality, i.e., the minimal cardinality of a subset of ${\omega^{\omega}}$ not having a given property. The critical cardinality for Menger is the dominating number $\mathfrak{d}$ and for Hurewicz it is the bounding number $\mathfrak{b}$. A set is totally imperfect if it does not contain a perfect subset, i.e., a homeomorphic copy of ${2^{\omega}}$. Known constructions of totally imperfect sets which are Menger or Hurewicz involve category-theoretic, measure-theoretic, combinatorial or forcing methods. However, each of these constructions provide a set satisfying ${\mathsf{S}_{1}(\Gamma,\mathrm{O})}$, which is another classical covering property well studied in the field [7, 19, 11]. A set $X$ is ${\mathsf{S}_{1}(\Gamma,\mathrm{O})}$ if for every sequence $\mathcal{U}_{0},\allowbreak\mathcal{U}_{1},\allowbreak\dotsc$ of open $\gamma$-covers of $X$, there are sets $U_{0}\in\mathcal{U}_{0},U_{1}\in\mathcal{U}_{1},\dotsc$ such that the family $\{\,U_{n}:n\in\omega\,\}$ covers $X$. The critical cardinality for ${\mathsf{S}_{1}(\Gamma,\mathrm{O})}$ is $\mathfrak{d}$ and the following implications hold [7].
@@
The only ZFC examples for separation Hurewicz or Menger from ${\mathsf{S}_{1}(\Gamma,\mathrm{O})}$ are not satisfactory, i.e., they are the Cantor space and, more generally, sets containing perfect subsets. Prior to this research, there hadn’t been totally imperfect examples, even under some additional assumptions beyond ZFC.
Using various set-theoretic assumptions which are weak portions of the Continuum Hypothesis, we construct totally imperfect Hurewicz or Menger sets which can be mapped continuously onto the Cantor cube. Each of already known strategies to construct totally imperfect sets with these properties have substantial limitations which prevent from reaching such a goal. These results have consequences for separation of considered properties. Another benefit is a solution of the below problem which was formulated independently by Nowik and Tsaban (in oral communication). By $\oplus$ we mean a standard group operation in ${2^{\omega}}$, i.e., for every $x,y\in{2^{\omega}}$ we have
@@
for all natural numbers $n$. For a set $A\subseteq{2^{\omega}}$ and an element $t\in{2^{\omega}}$, let $t\oplus A:=\{\,t\oplus a:a\in A\,\}$. A set $X\subseteq{2^{\omega}}$ is perfectly meager if for any perfect set $P\subseteq{2^{\omega}}$, the set $X\cap P$ is meager in the relative topology of $P$. By the result of Just–Miller–Scheepers–Szeptycki [7, Theorem 5.5] any totally imperfect Hurewicz subset of ${2^{\omega}}$ is perfectly meager. A set $X\subseteq{2^{\omega}}$ is perfectly meager in the transitive sense [12, 14] (PMT in short) if for any perfect set $P\subseteq{2^{\omega}}$, there is a $\sigma$-compact set $F\subseteq{2^{\omega}}$ containing $X$ such that the set $F\cap(t\oplus P)$ is meager in the relative topology of $t\oplus P$ for all $t\in{2^{\omega}}$.
Theorem 1.1 (Nowik [13, Theorem 1.]).
@@ 3. Menger sets
Let $\mathcal{U}_{0},\allowbreak\mathcal{U}_{1},\allowbreak\dotsc$ be a sequence of open covers of $X$. For each ordinal number $\alpha<\mathfrak{c}$ we have $w_{\alpha}\notin H_{\alpha}=\bigcap_{n\in\omega}\bigcup\mathcal{V}_{n}^{(\alpha)}$, and thus none of the sequences $\mathcal{V}_{0}^{(\alpha)},\mathcal{V}_{1}^{(\alpha)},\dotsc$ is a sequence of covers of $X$. Consequently, there is an ordinal number $\alpha<\mathfrak{c}$ such that the sequence $\mathcal{U}_{0},\allowbreak\mathcal{U}_{1},\allowbreak\dotsc$ is equal to $\mathcal{U}_{0}^{(\alpha)},\mathcal{U}_{1}^{(\alpha)},\dotsc$ By (ii) we have
@@ 4.2. Everywhere meager and everywhere null
A subset $A$ of a Polish space $X$ is universally meager [27, 28, 3, 4] if for any Borel isomorphism between $X$ and a Polish space $Y$, the image of $A$ under the map is meager. A set $X\subseteq{2^{\omega}}$ is everywhere meager [8] if for each set $a\in{[\omega]^{\omega}}$ the set $X_{a}:=\{\,x\upharpoonright a:x\in X\,\}$ is meager in $2^{a}$. Kraszewski proved that there is a universally meager set in ${2^{\omega}}$ which is not everywhere meager [8, Theorem 4.1]. We give a consistently negative answer to a weaker version of the following question.
Problem 4.1 (Kraszewski [8, Problem 1]).
@@
According to a characterization of universally meager sets due to Zakrzewski [27, Theorem 2.1], a subset of a Polish space is universally meager if and only if every Borel isomorphic image of this set into a Polish space is meager. It follows that universally meager sets are preserved by homeomorphisms. By the result of Pol and Zakrzewski [15, Theorem 3.1.3] consistently countably perfectly meager sets are not preserved by homeomorphisms. We do not know whether of everywhere meager sets are preserved by homeomorphisms.
4.3. The property $\mathsf{U}_{n}(\Gamma,\Gamma)$
A set $X$ satisfies $\mathsf{U}_{n}(\Gamma,\Gamma)$ if for any sequence $\mathcal{U}_{0},\allowbreak\mathcal{U}_{1},\allowbreak\dotsc$ of open $\gamma$-covers of $X$, there are sets $\mathcal{F}_{0}\subseteq\mathcal{U}_{0},\mathcal{F}_{1}\subseteq\mathcal{U}_{1},\dotsc$ such that $\left|\mathcal{F}_{n}\right|=n$ for all $n$’s and the family $\{\,\bigcup\mathcal{F}_{n}:n\in\omega\,\}$ is a $\gamma$-cover of $X$. This property was investigated by Tsaban [23] in the context of sets with the Hurewicz property and the property ${\mathsf{S}_{1}(\Gamma,\Gamma)}$. A set $X$ satisfies ${\mathsf{S}_{1}(\Gamma,\Gamma)}$ if for any sequence $\mathcal{U}_{0},\allowbreak\mathcal{U}_{1},\allowbreak\dotsc$ of open $\gamma$-covers of $X$, there are sets $U_{0}\in\mathcal{U}_{0},U_{1}\in\mathcal{U}_{1},\dotsc$ such that the family $\{\,U_{n}:n\in\mathbb{N}\,\}$ is a $\gamma$-cover of $X$. Each set satisfying ${\mathsf{S}_{1}(\Gamma,\mathrm{O})}$ is totally imperfect [23, Lemma 1.16]. Using a similar argument, we can show that any set satisfying $\mathsf{U}_{n}(\Gamma,\Gamma)$ is totally imperfect, too. A $\mathfrak{b}$-scale set is a subset of the Cantor space ${\operatorname{P}(\omega)}$ of the form $\{\,x_{\alpha}:\alpha<\mathfrak{b}\,\}\cup\mathrm{Fin}$, where $\{\,x_{\alpha}:\alpha<\mathfrak{b}\,\}$ is an unbounded subset of ${[\omega]^{\omega}}$ such that for every $\alpha<\beta<\mathfrak{b}$, we have $x_{\alpha}\leq^{*}x_{\beta}$. Such a set exists in ZFC and is an example of a Hurewicz totally imperfect set given by Bartoszyński–Shelah [5]. The only known example which distinguishes the properties ${\mathsf{S}_{1}(\Gamma,\Gamma)}$ and Hurewicz in the class of totally imperfect sets is a $\mathfrak{b}$-scale set [16] under $\mathfrak{b}=\mathfrak{c}$. By the results of Tsaban [23, Theorems 3.3 and 3.8] any $\mathfrak{b}$-scale set satisfies $\mathsf{U}_{n}(\Gamma,\Gamma)$, and thus any continuous image of a $\mathfrak{b}$-scale set into ${2^{\omega}}$ is totally imperfect. Tsaban asked the following question.
Problem 4.4 ([23, Problem 3.9]).
Is there a totally imperfect Hurewicz set which is not $\mathsf{U}_{n}(\Gamma,\Gamma)$?
Our example constructed in Theorem 2.1 provides a negative answer to the above question under $\operatorname{cov}(\mathcal{N})=\mathfrak{b}=\mathfrak{c}$.
@@ 5. Comments and open problems
Problem 5.1.
@@
Problem 5.3.

### v2

@@ Abstract.
The research of the first and second authors was funded by the National Science Center, Poland Weave-UNISONO call in the Weave programme Project: Set-theoretic aspects of topological selections 2021/03/Y/ST1/00122. The research of the third author was funded in whole by the Austrian Science Fund (FWF) [10.55776/I5930 and 10.55776/PAT5730424].
@@ 1. Introduction
Throughout this paper, we assume that a set $X$ with a given topological property is a subset of the Cantor space $2^{\omega}$} or its square. A set $X$ is Menger if for any sequence $\mathcal{U}_{0},\allowbreak\mathcal{U}_{1},\allowbreak\dotsc$ of open covers of $X$, there are finite sets $\mathcal{F}_{0}\subseteq\mathcal{U}_{0},\mathcal{F}_{1}\subseteq\mathcal{U}_{1},\dotsc$ such that the family $\{\,\bigcup\mathcal{F}_{n}:n\in\omega\,\}$ covers $X$. If in addition the family $\{\,\bigcup\mathcal{F}_{n}:n\in\omega\,\}$ is a $\gamma$-cover of $X$, i.e., for every $x\in X$, the set $\{\,n:x\in\bigcup\mathcal{F}_{n}\,\}$ is co-finite, then the set $X$ is Hurewicz. We have the following implications between considered properties.
$\sigma\text{-compact}\longrightarrow\text{Hurewicz}\longrightarrow\text{Menger}.$
A set is totally imperfect if it does not contain a perfect subset, i.e., a homeomorphic copy of $2^{\omega}$}. By the results of Bartoszyński–Shelah [5] and Tsaban–Zdomskyy [33], there are uniform constructions of totally imperfect sets in ZFC which show that none of the above implications are reversible in the realm of sets of reals. The above properties are central in the combinatorial covering properties theory and found applications in such areas as forcing [7], function spaces [26], Ramsey theory in algebra [31], combinatorics of discrete subspaces [1], hyperspaces with the Vietoris topology [11], products of Lindelöf spaces [2, 28] and products of paracompact spaces [27].
Each of the considered properties has a combinatorial characterization using continuous mappings into the Baire space $\omega^{\omega}$}. A set $X$ is Menger (Hurewicz) if and only if every continuous image of $X$ into $\omega^{\omega}$} is non-dominating (bounded) [23]. It follows that each of these properties has a critical cardinality, i.e., the minimal cardinality of a subset of $\omega^{\omega}$} not having a given property. The critical cardinality for Menger is the dominating number $\mathfrak{d}$ and for Hurewicz it is the bounding number $\mathfrak{b}$.
Many known constructions of totally imperfect sets which are Menger or Hurewicz involve category-theoretic, measure-theoretic, combinatorial or forcing methods. However, each of these constructions provide a set satisfying $\mathsf{S}_{1}(\Gamma,\mathrm{O})$}, which is another classical covering property well studied in the field [8, 24, 14]. A set $X$ is $\mathsf{S}_{1}(\Gamma,\mathrm{O})$} if for every sequence $\mathcal{U}_{0},\allowbreak\mathcal{U}_{1},\allowbreak\dotsc$ of open $\gamma$-covers of $X$, there are sets $U_{0}\in\mathcal{U}_{0},U_{1}\in\mathcal{U}_{1},\dotsc$ such that the family $\{\,U_{n}:n\in\omega\,\}$ covers $X$. The critical cardinality for $\mathsf{S}_{1}(\Gamma,\mathrm{O})$} is $\mathfrak{d}$ and the following implications hold [8].
@@
The only ZFC examples for separation Hurewicz or Menger from $\mathsf{S}_{1}(\Gamma,\mathrm{O})$} are not satisfactory, i.e., they are the Cantor space and, more generally, sets containing perfect subsets. After submitting the paper, Zakrzewski shared with us earlier hand notes of Pol with similar constructions as presented in our Theorems 2.1 and 3.1, where in particular we construct a totally imperfect Hurewicz set and totally imperfect Menger set, using additional set-theoretic assumptions beyond ZFC. Professor Roman Pol also pointed out to us that our Theorem 3.1 can be deduced from papers [18, 19] written in a different context and using different techniques.
A set $X\subseteq{2^{\omega}}$ is perfectly meager if for any perfect set $P\subseteq{2^{\omega}}$, the set $X\cap P$ is meager in the relative topology of $P$. By the result of Just–Miller–Scheepers–Szeptycki [8, Theorem 5.5] any totally imperfect Hurewicz subset of $2^{\omega}$} is perfectly meager. It follows from the result of Bartoszyński–Shelah [5] that in ZFC there is a perfectly meager set which is not Hurewicz. Let $\operatorname{cof}(\mathcal{M})$ be the minimal cardinality of a family of meager sets in $2^{\omega}$} such that each meager set in $2^{\omega}$} is contained in a memebr of the family. Using techinques from the work of Szewczak–Wiśniewski [29, Proposition 2.9.], it can be shown that assuming the equality $\mathfrak{d}=\operatorname{cof}(\mathcal{M})$, there is a totally imperfect Menger set which is not perfectly meager.
Using various set-theoretic assumptions which are weak portions of the Continuum Hypothesis, we construct totally imperfect Hurewicz or Menger sets which can be mapped continuously onto the Cantor cube. Most of already known strategies to construct totally imperfect sets with these properties have substantial limitations which prevent from reaching such a goal. These results have consequences for separation of considered properties. Another benefit is a solution of the below problem which was formulated independently by Nowik and Tsaban (in oral communication). By $\oplus$ we mean a standard group operation in $2^{\omega}$}, i.e., for every $x,y\in{2^{\omega}}$ we have
@@
for all natural numbers $n$. For a set $A\subseteq{2^{\omega}}$ and an element $t\in{2^{\omega}}$, let $t\oplus A:=\{\,t\oplus a:a\in A\,\}$. A set $X\subseteq{2^{\omega}}$ is perfectly meager in the transitive sense [15, 17] (PMT in short) if for any perfect set $P\subseteq{2^{\omega}}$, there is a $\sigma$-compact set $F\subseteq{2^{\omega}}$ containing $X$ such that the set $F\cap(t\oplus P)$ is meager in the relative topology of $t\oplus P$ for all $t\in{2^{\omega}}$.
Theorem 1.1 (Nowik [16, Theorem 1.]).
@@ 3. Menger sets
Let $\mathcal{U}_{0},\allowbreak\mathcal{U}_{1},\allowbreak\dotsc$ be a sequence of open covers of $X$. For each ordinal number $\alpha<\mathfrak{c}$ we have $w_{\alpha}\notin H_{\alpha}=\bigcap_{n\in\omega}\bigcup\mathcal{V}_{n}^{(\alpha)}$, and thus none of the sequences $\mathcal{V}_{0}^{(\alpha)},\mathcal{V}_{1}^{(\alpha)},\dotsc$ is a sequence of covers of $X$. Consequently, there is an ordinal number $\alpha<\mathfrak{c}$ such that the sequence $\mathcal{U}_{0},\allowbreak\mathcal{U}_{1},\allowbreak\dotsc$ is equal to $\mathcal{U}_{0}^{(\alpha)},\mathcal{U}_{1}^{(\alpha)},\dotsc$ . By (ii) we have
@@ 4.2. Everywhere meager and everywhere null
A subset $A$ of a Polish space $X$ is universally meager [35, 36, 3, 4] if for any Borel isomorphism between $X$ and a Polish space $Y$, the image of $A$ under the map is meager. A set $X\subseteq{2^{\omega}}$ is everywhere meager [10] if for each set $a\in{[\omega]^{\omega}}$ the set $X_{a}:=\{\,x\upharpoonright a:x\in X\,\}$ is meager in $2^{a}$. Kraszewski proved that there is a universally meager set in $2^{\omega}$} which is not everywhere meager [10, Theorem 4.1]. We give a consistently negative answer to a weaker version of the following question.
Problem 4.1 (Kraszewski [10, Problem 1]).
@@
According to a characterization of universally meager sets due to Zakrzewski [35, Theorem 2.1], a subset of a Polish space is universally meager if and only if every Borel isomorphic image of this set into a Polish space is meager. It follows that universally meager sets are preserved by homeomorphisms. By the result of Pol and Zakrzewski [20, Theorem 3.1.3] consistently countably perfectly meager sets are not preserved by homeomorphisms. We do not know whether of everywhere meager sets are preserved by homeomorphisms.
@@ 5. Comments and open problems
5.1. Groupability and considered properties
Considered here properties as Hurewicz and Menger have also equivalent formulations using another types of open covers of spaces. Let $\mathcal{U}$ be a countable cover of a set $X$. The family $\mathcal{U}$ is an $\omega$-cover of $X$, if $X\notin U$ and each finite subset of $U$ is contained in a set from $\mathcal{U}$. The family $\mathcal{U}$ is groupable, if there is a partition $\{\,\mathcal{F}_{n}:n\in\omega\,\}$ of $\mathcal{U}$ into pairwise disjoint finite sets such that the family $\{\,\bigcup\mathcal{F}_{n}:n\in\omega\,\}$ is a $\gamma$-cover of $X$. A set $X$ satsifies $\mathsf{S}_{\mathrm{fin}}(\Omega,\mathrm{O})$ if for any sequence $\mathcal{U}_{0},\allowbreak\mathcal{U}_{1},\allowbreak\dotsc$ of open $\omega$-covers of $X$, there are finite sets $\mathcal{F}_{0}\subseteq\mathcal{U}_{0},\mathcal{F}_{1}\subseteq\mathcal{U}_{1},\dotsc$ such that the family $\bigcup_{n\in\omega}\mathcal{F}_{n}$ is a cover of $X$. A set $X$ satsifies $\mathsf{S}_{\mathrm{fin}}(\Omega,\mathrm{O}^{\mathrm{gp}})$ if for any sequence $\mathcal{U}_{0},\allowbreak\mathcal{U}_{1},\allowbreak\dotsc$ of open $\omega$-covers of $X$, there are finite sets $\mathcal{F}_{0}\subseteq\mathcal{U}_{0},\mathcal{F}_{1}\subseteq\mathcal{U}_{1},\dotsc$ such that the family $\bigcup_{n\in\omega}\mathcal{F}_{n}$ is a groupable cover of $X$. By a result of Scheepers [25] the Menger property is equivalent to $\mathsf{S}_{\mathrm{fin}}(\Omega,\mathrm{O})$ and by the result of Kocinac–Scheepers [9, Theorem 14] the Hurewicz property is equivalent to $\mathsf{S}_{\mathrm{fin}}(\Omega,\mathrm{O}^{\mathrm{gp}})$. Our paper provides another than already known examples of $\mathsf{S}_{\mathrm{fin}}(\Omega,\mathrm{O})$ totally imperfect sets which are not $\mathsf{S}_{\mathrm{fin}}(\Omega,\mathrm{O}^{\mathrm{gp}})$. Using dichotomic proof, Chaber–Pol [6] proved that there is such a set in ZFC. A unfiorm ZFC construction of such a set was provied by Tsaban–Zdomskyy [33].
5.2. The property $\mathsf{U}_{\mathrm{id}}(\Gamma,\Gamma)$
A set $X$ satisfies $\mathsf{U}_{\mathrm{id}}(\Gamma,\Gamma)$, if for any sequence $\mathcal{U}_{0},\allowbreak\mathcal{U}_{1},\allowbreak\dotsc$ of open $\gamma$-covers of $X$, there are sets $\mathcal{F}_{0}\subseteq\mathcal{U}_{0},\mathcal{F}_{1}\subseteq\mathcal{U}_{1},\dotsc$ such that $\left|\mathcal{F}_{n}\right|=n$ for all $k$’s and the family $\{\,\bigcup\mathcal{F}_{n}:n\in\omega\,\}$ is a $\gamma$-cover of $X$. Let $k$ be a positive natural number. A set $X$ satisfies $\mathsf{U}_{k}(\Gamma,\Gamma)$, if for any sequence $\mathcal{U}_{0},\allowbreak\mathcal{U}_{1},\allowbreak\dotsc$ of open $\gamma$-covers of $X$, there are sets $\mathcal{F}_{0}\subseteq\mathcal{U}_{0},\mathcal{F}_{1}\subseteq\mathcal{U}_{1},\dotsc$ such that $\left|\mathcal{F}_{n}\right|=k$ for all $n$’s and the family $\{\,\bigcup\mathcal{F}_{n}:n\in\omega\,\}$ is a $\gamma$-cover of $X$. This property was investigated by Tsaban [30] in the context of sets with the Hurewicz property and the property $\mathsf{S}_{1}(\Gamma,\Gamma)$}. A set $X$ satisfies $\mathsf{S}_{1}(\Gamma,\Gamma)$} if for any sequence $\mathcal{U}_{0},\allowbreak\mathcal{U}_{1},\allowbreak\dotsc$ of open $\gamma$-covers of $X$, there are sets $U_{0}\in\mathcal{U}_{0},U_{1}\in\mathcal{U}_{1},\dotsc$ such that the family $\{\,U_{n}:n\in\omega\,\}$ is a $\gamma$-cover of $X$. We have the following implications
[…]
Tsaban asked the following question.
Problem 5.1 ([30, Problem 3.9]).
Is there a totally imperfect Hurewicz set which is not $\mathsf{U}_{\mathrm{id}}(\Gamma,\Gamma)$?
Our example constructed in Theorem 2.1 provides a negative answer to the above question under $\operatorname{cov}(\mathcal{N})=\mathfrak{b}=\mathfrak{c}$. Assume that $\mathfrak{b}=fc$. In a very recent paper [32], using different techniques Tsaban proved that there is a Hurewicz set satisfying $\mathsf{S}_{1}(\Gamma,\mathrm{O})$} (and thus, it is totally imperfect) which is not $\mathsf{U}_{\mathrm{id}}(\Gamma,\Gamma)$  [32, Theorem 10.6]. Tsbaban also proved that there is a set satisfying $\mathsf{U}_{\mathrm{id}}(\Gamma,\Gamma)$ but not $\mathsf{U}_{k}(\Gamma,\Gamma)$ for any positive natural number $k$ [32, Theorem 9.9] and also that for each $k$ there is a set satisfying $\mathsf{U}_{k+1}(\Gamma,\Gamma)$ but not $\mathsf{U}_{k}(\Gamma,\Gamma)$[32, Theorem 10.3] There are also another results of Liu–He–Zhang in a similar spirit, if we assume that the Continuum Hypothesis holds [12].
5.3. Separation of properties
Problem 5.2.
Is there in ZFC a totally imperfect Menger set which is not perfectly meager? Is there in ZFC a perfectly meager set which is not Menger [^1] In a recent work of V. Haberl with the first and third named authors it is shown that every perfectly meager set has size $\omega_{1}$ in the Miller model, and hence all such sets are Menger there.?
Problem 5.3.
5.4. Around PMT sets
Problem 5.4.
@@ 5.4. Around PMT sets
Acknowledgments
We would like to thank to the anonymous referee for his/her suggestions and comments which help us to increase substantially the quality of this paper. We also would like to thank to Professor Piotr Zakrzewski and Professor Roman Pol for fruitful discussions about obtained in the paper results.